# Adapting Pre-Trained Vision Models for Novel Instance Detection and Segmentation

### v2

@@ 8 Unseen Detection of BOP Benchmark
NIDS-Net + WA (Ours) + $s_{\text{appe}}$ 45.7 49.3 48.6 25.7 37.9 58.7 62.1 46.9
9 More Qualitative Results
9.1 Adapter
To facilitate a comparison between the CLIP-Adapter and our Weight Adapter, we present a visual illustration in Figure 7. It is evident that the CLIP-Adapter alters the feature space and spoils the embeddings of background objects due to overfitting. In contrast, our Weight Adapter delivers robust embeddings within the original feature space.
9.2 Detection
We display the visual outcomes of our methodology employing the weight adapter on the LMO and YCB-V datasets in Figure 8. The gap between synthetic and real images results in some instances of detection failure. For example, some LM-O instances are not found with our method.
9.3 Segmentation
We present the visual outcomes of our approach using the weight adapter on the BoP datasets in Figures 9, 10, 11, and 12. These images demonstrate the effectiveness of our approach in cluttered environments. In some cases of T-LESS and IC-BIN datasets, Grounding DINO generates large bounding boxes which include multiple objects, causing under-segmentation. Furthermore, in IC-BIN and HB datasets, some heavily occluded objects with low confidence scores are overlooked by our method.
Figure 7: Comparison of different adapters on a hard scene image from the high-resolution dataset. Red arrows denote background objects that are erroneously classified as targets.

### v3

@@ VIII Unseen Detection of BOP Benchmark
NIDS-Net + WA + $s_{\text{appe}}$ (Ours) 45.7 49.3 48.6 25.7 37.9 58.7 62.1 46.9
IX More Qualitative Results
IX-A Adapter
To facilitate a comparison between the CLIP-Adapter and our Weight Adapter, we present a visual illustration in Figure 9. It is evident that the CLIP-Adapter alters the feature space and spoils the embeddings of non-target objects due to overfitting. In contrast, our Weight Adapter delivers robust embeddings within the original feature space.
IX-B Detection
We display the visual outcomes of our methodology with the weight adapter on the LMO and YCB-V datasets in Figure 11. The gap between synthetic and real images results in some instances of detection failure. For example, some LM-O instances are not found with our method. The examples of the high-resolution dataset are presented in Fig. 10.
IX-C Segmentation
We present the visual results of our approach using the weight adapter on the BoP datasets in Figures 12, 13, 14, and 15. These images demonstrate the effectiveness of our approach in cluttered environments. In some cases of T-LESS and IC-BIN datasets, Grounding DINO generates large bounding boxes which include multiple objects, causing under-segmentation. Furthermore, in IC-BIN and HB datasets, some heavily occluded objects with low confidence scores are overlooked by our method.
Figure 9: Comparison of different adapters on a hard scene image from the high-resolution dataset. Red arrows denote non-target objects that are erroneously classified as targets.
Figure 10: Visual examples of our results using the Weight Adapter on the high-resolution dataset. Our approach detects specific object instances in cluttered scenes according to their real template images.